OpenStudy (anonymous):

Identify the sequence as arithmetic, geometric, or neither. Explain your answer. 1.6, 0.8, 0.4, 0.2, . . .

OpenStudy (anonymous):

idk what do u think?

OpenStudy (anonymous):

arithmetic?

OpenStudy (anonymous):

huh..onestly idk it either

OpenStudy (anonymous):

okay then.....

OpenStudy (anonymous):

hope u get the answer anyways

OpenStudy (anonymous):

@mmmmmmmmmm123 ?!?!

myininaya (myininaya):

is 1.6-0.8 the same as 0.8-0.4

myininaya (myininaya):

are both of those differences the same?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

yes

myininaya (myininaya):

so this is definitely not a arithmetic sequence since there is no common difference amomgst the consecutive numbers in your set

myininaya (myininaya):

now let's look at ratios

myininaya (myininaya):

is 1.6/0.8 the same as 0.8/0.4?

OpenStudy (anonymous):

no?

myininaya (myininaya):

?

myininaya (myininaya):

16/8 isn't the same as 8/4?

OpenStudy (anonymous):

yess sorry it is

myininaya (myininaya):

and 16/8=8/4=4/2 all of these ratios equal 2

myininaya (myininaya):

so you have a common ratio

myininaya (myininaya):

which means you have what kinda sequence?

OpenStudy (anonymous):

geometric

myininaya (myininaya):

common difference goes with arithmetic common ratio goes with geometric yes you have a geometric sequence

OpenStudy (anonymous):

okayyy thanks!!!!

myininaya (myininaya):

np